The college student senate is sponsoring a spring break Caribbean cruise raffle. The proceeds are to be donated to the Samaritan Center for the Homeless. A local travel agency donated the cruise, valued at $2000. The students sold 2867 raffle tickets at $5 per ticket.

(a) Kevin bought twenty-five tickets. What is the probability that Kevin will win the spring break cruise to the Caribbean? (Round your answer to five decimal places.) What is the probability that Kevin will not win the cruise? (Round your answer to five decimal places.)

(b) Expected earnings can be found by multiplying the value of the cruise by the probability that Kevin will win. What are Kevin's expected earnings? (Round your answer to two decimal places.)

$

(c) Is this more or less than the amount Kevin paid for the twenty-five tickets?

(d) How much did Kevin effectively contribute to the Samaritan Center for the Homeless? (Round your answer to two decimal places.)

$
Answer #1
a)P(kevin wins) =25/2867= 0.00872
P(kevin not win) =1-0.00872= 0.99128

b)

expected earnings =2000*0.00872= 17.44

c

less

d)

effective contribution=25*5-17.44= 107.56